The length of the skid mark of an automobile sliding to a stop is directly proportional to the square of the velocity. It was found that a car traveling 60 miles per hour left a skid mark 172 feet in length when attempting to stop. Find the length of the skid marks of a car traveling 50 miles per hour.
Answer by RAY100(1637)

Let L= skid length in ft,,,,v= veolcity in MPH
.
L=kv^2
.
subst data given,,,,,172 = k (60)^2,,,,,,,k= 172/60^2 = .0477
.
proportion becomes,,,L= .0477 v^2
.
for v=50,,,,,L = .0477 * 50^2 = 119.44 ft
.
.
checking,,,, 119.44/50^2 = .0477,,,,ok